OpenStudy (anonymous):

I need to get 6x + 3y =-9 x into slope intercept form I got y= -6/-3 1 is this right

OpenStudy (anonymous):

6x +3y =-9 -6x+6x+3y = -9 -6x 3y/3=-9 -6x/3 y =-9/3 -6x/3 y = -6x/3 -9/3 y = -2x -9/3 y intercept form

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

But slope comes first

OpenStudy (anonymous):

the slope is -2. i don't understand "slope comes first"

OpenStudy (anonymous):

I thought the slope had -9/3

OpenStudy (anonymous):

Was

OpenStudy (anonymous):

y = mx + b m is the slope b is the y intercept....where the line crosses the y axis

OpenStudy (anonymous):

Ok thank you will you check this 5x - 2y =-2 -5x = -2 -2y=-2+5x Y = 1 5x/-2

OpenStudy (anonymous):

it should be 5x -2y =-2 -5x +5x -2y = -2 -5x -2y = -2 -5x -2/-2 y = -2/-2 -5/-2 x y = 1 + 5/2 x rearrange this to y = 5/2x + 1